# Next Generation Equation-Free Multiscale Modelling of Crowd Dynamics via Machine Learning

### v2

@@ Abstract
Bridging the microscopic/individual and macroscopic/emergent modelling scales in crowd dynamics constitutes an important, open challenge for systematic numerical analysis, optimization, and control. Here, we propose a combined manifold and machine learning approach to learn the discrete evolution operator for the emergent/collective crowd dynamics in latent spaces from high-fidelity individual/agent-based simulations. The proposed framework builds upon our previous works on the development of next-generation Equation-free algorithms for dealing with the “curse of dimensionality” when learning surrogate models for high-dimensional and multiscale systems. The proposed approach is a four-stage one, explicitly conserving the mass of the reconstructed dynamics in the high-dimensional space. In the first step, we derive continuous macroscopic fields (densities) from discrete microscopic data (pedestrians’ positions) using Kernel Density Estimation. In the second step, based on manifold learning, we construct a map from the macroscopic ambient space into an appropriate latent space parametrized by a few coordinates based on Proper-Orthogonal Decomposition (POD) of the corresponding density distributions. The third step involves learning reduced-order surrogate models in the latent space using machine learning techniques, particularly Long Short-Term Memory (LSTMs) networks and Multivariate Autoregressive models (MVARs). Finally, we reconstruct the crowd dynamics in the high-dimensional space in terms of macroscopic density profiles. We demonstrate that the POD reconstruction of the density distribution via Singular Value Decomposition conserves the mass. With this “embed $\rightarrow$ learn in latent space $\rightarrow$ lift back to the high-dimensional space” pipeline, we create an effective solution operator of the unavailable (at the macroscopic scale) Partial Differential Equation for the evolution of the density distribution. For our illustrations, we use the Social Force Model to generate data in a corridor with an obstacle, imposing periodic boundary conditions. The numerical results demonstrate high accuracy, robustness, and generalizability, thus allowing for fast and accurate modelling/simulation of crowd dynamics from agent-based simulations. Notably, linear MVAR models surpass nonlinear LSTMs in predictive accuracy, while also offering significantly lower complexity and greater interpretability.
@@ 1 Introduction
An important open challenge in the modelling of human crowd dynamics lies in bridging across the different modelling scales. Typically, models are formulated at three distinct levels: the microscopic (individual), mesoscopic (kinetic), and macroscopic (hydrodynamic) scales. Microscopic models describe the detailed behavior of individual pedestrians using agent-based approaches, such as the Social Force Model (SFM) [1] and its variants (see e.g. [2, 3]), cellular automata [4, 5, 6], etc. The mesoscopic scale serves as an intermediary between the microscopic and macroscopic approaches [7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. At the macroscopic scale, modelling describes the emergent/collective behavior of crowds. For instance, in the seminal work of Hughes [17], the author models crowd dynamics by coupling a conservation law with an eikonal equation to represent how individuals choose paths based on a potential field. Moreover, in [18] the authors proposed nonlocal conservation laws for traffic and crowd dynamics, where the flux depends not only on local density but also on averaged information from the surrounding region. In [11], a framework for macroscopic crowd models was presented based on balance laws and kinetic theory, emphasizing how individual-level interactions give rise to complex collective behavior. More recently, [16] highlights new trends such as multiscale coupling, behavioral heterogeneity, and nonlocal interaction kernels.
@@
It is important to note, however, that despite the long-standing tradition of modeling crowd dynamics through (hydrodynamic) PDEs, such formulations inherently rely on an “infinite-size” assumption. In realistic settings involving crowds of moderate or finite-size, this assumption often breaks down. Finite-size effects, such as spatial heterogeneity, individual interactions, and boundary influences, can dominate the dynamics, rendering continuum PDEs insufficient for accurate, quantitative predictions of real-world crowd behavior. Here, inspired by the Equation-free multiscale framework [39], and building on our previous recent works, on what we call next-generation Equation-free algorithms based on both manifold learning and machine learning [62, 42, 63, 27, 64, 47], we present a computational framework, to learn the discrete evolution operator of the emergent dynamics in latent spaces from high-fidelity microscopic simulations of crowd dynamics. The proposed framework bridges the microscopic and macroscopic modelling scales, while explicitly preserving the underlying conservation law of mass in the ambient space. The methodology is deployed in four main stages. First, we map discrete microscopic distributions to continuous macroscopic representations (i.e., density fields). Second, we project the macroscopic fields into a low-dimensional latent space using Proper Orthogonal Decomposition (POD). Third, we learn surrogate reduced-order models (ROMs) in this latent space using discrete autoregressive maps: both linear Multivariate Autoregressive models (MVARs) and nonlinear LSTM models. These autoregressive models with lags implicitly implement a delay coordinate embedding, aligning with Takens’ theorem [65] for reconstructing the phase-space structure (here in the latent space). The concept of the approach (i.e., using POD embeddings and neural networks taking as inputs delayed time series for learning the dynamics of PDEs in latent spaces) can be traced back to the early ’90s [66, 67] (see also results and discussion in [67, 62, 68, 69, 63]). In crowd dynamics, behavioral characteristics such as pedestrian target paths and decision-making, or path optimization constitute the “unobservable” states, driving the observable positions/velocities and therefore the density field of the crowd; delay embeddings (as used implicitly in autoregressive models) have the potential, due to the Takens’ theorem, to reconstruct the effective phase space containing these (unobservable) latent dynamics. Finally, we lift the learned latent-space dynamics back to the high-dimensional macroscopic space via the POD basis. Crucially, this configuration ensures explicit mass conservation—a key physical constraint for any effective continuum description. For our numerical experiments, we employ the Social Force Model (SFM) [70] to simulate pedestrian flow within a rectangular corridor containing an obstacle, under periodic boundary conditions. The performance and generalization ability of the proposed framework are assessed on unseen initial conditions.
@@ 2.1 Problem Statement
where $\Phi_{\Delta t}:\mathbb{R}^{M}\to\mathbb{R}^{M}$ is the evolution operator at the high-dimensional space. In this work, we learn the discrete evolution operator $F_{\delta t}$ using autoregressive ROMs, thus effectively learning $\Phi_{\Delta t}$ via the restriction and lifting operators, as:
$\mathbf{u}(t+\Delta t)=\mathcal{L}\Bigl(F_{\delta t}\big(\mathcal{R}(\mathbf{u}(t)),\ldots,\mathcal{R}(\mathbf{u}(t-w\delta t))\big)\Bigl),$ (5)
where the learned ROMs account for $w$ delayed embeddings $\mathcal{R}(\mathbf{u}(t-j\delta t))$ for $j=1,\ldots,w$.
@@ 2.2.1 From discrete pedestrian positions to continuous density fields
The initial step in our framework involves constructing a continuous density field from the discrete positions of pedestrians. Given microscopic observations, obtained by agent-based simulators in the form of Eq. (1, 2) or experimental data, macroscopic continuous quantities are typically derived via local averaging at each point of the spatial domain where the crowd moves. Here, we start from $N$ pairs of pedestrian positions $\{\mathbf{x}_{i}(t)\}_{i=1}^{N}\subset\Omega$ in a rectangular domain $\Omega=[a,b]\times[c,d]\subset\mathbb{R}^{2}$, as obtained by the SFM microscopic simulator, detailed in Appendix A. We next employ KDE [84] to estimate the continuous density fields along points on the spatial grid $\mathbf{x}\in\Omega$, as:
@@
where $K_{h}$ is a smooth non-negative kernel function satisfying $\int_{\Omega}K(\mathbf{x})\,d\mathbf{x}=1$. Using a standard uniform spatial discretization $\delta x=(b-a)/n_{x}$, $\delta y=(c-d)/n_{y}$ with $n_{x}$$n_{y}$ grid points along the xx/$y$-axis of the domain $\Omega$, the resulting from Eq. (6) density fields are computed as $\rho(\mathbf{x},t)=[\rho((x_{1},y_{1}),t),\ldots,\rho((x_{n_{x}},y_{n_{y}}),t)]^{\top}\in\mathbb{R}^{n_{x}\times n_{y}}$, where $\rho((x_{i},y_{j}),t)$ corresponds to the density computed at the grid point $(x_{i},y_{j})$, for $i=1,\dots,n_{x}$, $j=1,\dots,n_{y}$. Then, the total mass in the domain is given by $S(t)=\sum_{i=1}^{n_{x}}\sum_{j=1}^{n_{y}}\rho((x_{i},y_{j}),t)\,\delta x\,\delta y$. While KDE can be used for extracting continuous macroscopic fields, such as momentum and energy, from any microscopic discrete quantities, here we only construct the density fields in Eq. (6). Additional details for the discretization of the computational domain $\Omega$, the choice of kernel $K_{h}(\cdot)$, the consideration of obstacles in $\Omega$ (where a binary mask function is employed to consider zero density at the grid points of the obstacle), and the treatment of periodic boundary conditions are provided in Appendix C.
@@ 2.2.2 The Restriction and Lifting Operators
where $n_{c}=n_{x}\times n_{y}$. To construct the restriction operator, we apply POD [83] to a dataset consisting of macroscopic density “solutions” in time, obtained by microscopic simulations via Eq. (6), for different initial conditions; see Appendix B for details. Denoting the densities by $\rho^{(c)}(\mathbf{x},t_{0}+k\delta t)\equiv\rho^{(c)}(\mathbf{x},t_{k})\in\mathbb{R}^{n_{c}}$, where $\delta t$ is a microscopic inner simulation step/observation sampling time of an experiment/real data, $t_{0}$ is the initial time, and the superscript $(c)$ denotes the case $c=1,\ldots,C$ of the different initial conditions considered, the training dataset generated by microscopic runs is represented by the matrix $X\in\mathbb{R}^{n_{c}\times(C\cdot K)}$:
$X=\Bigl\{\rho^{(c)}(\mathbf{x},t_{k})\,:\,k=0,\ldots,K,\,c=1,\ldots,C\Bigr\}.$ (8)
@@
where $U\in\mathbb{R}^{n_{c}\times n_{c}}$ is the matrix collecting the left-singular vectors, $\Sigma\in\mathbb{R}^{n_{c}\times n_{t}}$ is the matrix collecting the singular vectors, $V\in\mathbb{R}^{n_{t}\times n_{t}}$ is the matrix collecting the right-singular vectors, $I_{n_{t}}$ is the $n_{t}$-dim. unity matrix and $\mathbf{1}_{n_{t}}$ is the $n_{t}$-dim. unit column vector. The latent space is spanned by the first $d$ left-singular vectors forming $U_{d}$. Thus, the projection of $X$ on the latent space provides the low-dimensional embedding:
$Y_{d}=U_{d}^{\top}\bar{X},\,\,Y_{d}\in\mathbb{R}^{d\times n_{t}}$ (11)
@@
In summary, given a density distribution $\rho(\mathbf{x},t)$ obtained via Eq. (6), we obtain the normalized density along the $n_{c}$ grid points of $\Omega$ as:
$x(t)=\dfrac{1}{S(t)}[\rho((x_{1},y_{1}),t),\ldots,\rho((x_{n_{x}},y_{n_{y}}),t)]^{\top}\in\mathbb{R}^{n_{c}},$ (19)

### v3

# arXiv:2508.03926v3[cs.LG] 28 Feb 2026

@@ Abstract
Bridging the microscopic/individual and macroscopic/emergent modelling scales in crowd dynamics constitutes an important, open challenge for systematic numerical analysis, optimization, and control. Here, we propose a manifold-informed machine learning approach to learn the discrete evolution operator for the emergent/collective crowd dynamics in latent spaces from high-fidelity individual/agent-based simulations. The proposed framework builds upon our previous works on the development of next-generation Equation-free algorithms for dealing with the “curse of dimensionality” when learning surrogate models for high-dimensional and multiscale systems. The proposed approach is a four-stage one, explicitly conserving the mass of the reconstructed dynamics in the high-dimensional space. In the first step, we derive continuous macroscopic fields (densities) from discrete microscopic data (pedestrians’ positions) using Kernel Density Estimation. In the second step, based on manifold learning, we construct a map from the density-field space into an appropriate latent space parametrized by a few coordinates based on Proper-Orthogonal Decomposition (POD) of the corresponding density distributions. The third step involves learning reduced-order surrogate models in the latent space using machine learning techniques, particularly Long Short-Term Memory networks and Multivariate Autoregressive models. Finally, we reconstruct the crowd dynamics in the high-dimensional space with POD, demonstrating that the POD reconstruction conserves the mass. Thus, with this “embed $\rightarrow$ learn in latent space $\rightarrow$ lift back to the high-dimensional space” pipeline, we create an effective solution operator of the unavailable (at the macroscopic scale) Partial Differential Equation for the evolution of the density distribution. For our illustrations, we used the Social Force Model to generate data in a corridor with an obstacle, imposing periodic boundary conditions in two scenarios: (i) an unidirectional flow, and (ii) a counterflow. The numerical results demonstrate high accuracy, robustness, and generalizability, thus allowing for fast and accurate modelling/simulation of crowd dynamics from agent-based simulations.
@@ 1 Introduction
An important open challenge in the modelling of human crowd dynamics lies in bridging across the different modelling scales. Typically, models are formulated at three distinct levels: the microscopic (individual), mesoscopic (kinetic), and macroscopic (hydrodynamic) scales. Microscopic models describe the detailed behavior of individual pedestrians using agent-based approaches, such as the Social Force Model (SFM) [62, 61] and its variants (see e.g. [32, 111]), cellular automata [87, 21, 41], etc. The mesoscopic scale serves as an intermediary between the microscopic and macroscopic approaches [11, 15, 35, 13, 16, 43, 72, 66, 10, 19]. At the macroscopic scale, modelling describes the emergent/collective behavior of crowds. For instance, in the seminal work of Hughes [68], the author models crowd dynamics by coupling a conservation law with an eikonal equation to represent how individuals choose paths based on a potential field. Moreover, in [31] the authors proposed nonlocal conservation laws for traffic and crowd dynamics, where the flux depends not only on local density but also on averaged information from the surrounding region. In [16], a framework for macroscopic crowd models was presented based on balance laws and kinetic theory, emphasizing how individual-level interactions give rise to complex collective behavior. More recently, [19] highlights new trends such as multiscale coupling, behavioral heterogeneity, and nonlocal interaction kernels.
@@
It is important to note, however, that despite the long-standing tradition of modeling crowd dynamics through (hydrodynamic) PDEs, such formulations inherently rely on an “infinite-size” assumption. In realistic settings involving crowds of moderate or finite-size, this assumption often breaks down. Finite-size effects, such as spatial heterogeneity, individual interactions, and boundary influences, can dominate the dynamics, rendering continuum PDEs insufficient for accurate, quantitative predictions of real-world crowd behavior.
@@
Here, inspired by the Equation-free multiscale framework [51], and building on our previous recent works, on what we call next-generation Equation-free algorithms based on both manifold learning and machine learning [46, 93, 47, 37, 91, 49], we present a computational framework, to learn the discrete evolution operator of the emergent dynamics in latent spaces from high-fidelity microscopic simulations of crowd dynamics. The proposed framework bridges the microscopic and macroscopic modelling scales, while explicitly preserving the underlying conservation law of mass in the ambient space. The methodology is deployed in four main stages. First, we map discrete microscopic distributions to continuous macroscopic representations (i.e., density fields). Second, we project the macroscopic fields into a low-dimensional latent space using Proper Orthogonal Decomposition (POD). Third, we learn surrogate reduced-order models (ROMs) in this latent space using discrete autoregressive maps: both linear Multivariate Autoregressive models (MVARs) and nonlinear LSTM models. These autoregressive models with lags implicitly implement a delay coordinate embedding, aligning with Takens’/Whitney’s embedding theorems [106, 99] for reconstructing the phase-space structure (here in the latent space). The concept of the approach (i.e., using POD embeddings and neural networks taking as inputs delayed time series for learning the dynamics of PDEs in latent spaces) can be traced back to the early ’90s [75, 102] (see also results and discussion in [91, 49, 102, 71, 9]). In crowd dynamics, behavioral characteristics such as pedestrian target paths and decision-making, or path optimization constitute the “unobservable” states, driving the observable positions/velocities and therefore the density field of the crowd; delay embeddings (as used implicitly in autoregressive models) have the potential, due to the Takens’/Whitney’s embedding theorems [106, 99], to reconstruct the effective phase space containing these (unobservable) latent dynamics. Finally, we lift on-demand the learned latent-space dynamics back to the high-dimensional macroscopic space via the POD basis. Crucially, this configuration ensures explicit mass conservation-—a key physical constraint for any effective continuum description. For our numerical experiments, we employ the SFM [62, 61] to simulate pedestrian flow within a rectangular corridor containing an obstacle, under periodic boundary conditions. We particularly consider a unidirectional pedestrian flow and a more challenging, dense counterflow of two pedestrian groups. The performance and generalization ability of the proposed framework are assessed on unseen initial conditions.
@@ 2.1 Problem Statement
where $\Phi_{{\delta}t}:\mathbb{R}^{M}\to\mathbb{R}^{M}$ is the evolution operator at the high-dimensional space. In this work, we learn the discrete evolution operator $F_{\delta t}$ using autoregressive ROMs, thus effectively learning $\Phi_{{\delta}t}$ as:
$\mathbf{u}(t+\delta t)=\mathcal{L}\left(\mathbf{y}_{d}(t+\delta t)\right),\quad\mathbf{y}_{d}(t+\delta t)=F_{\delta t}\big(\mathbf{y}_{d}(t),\ldots,\mathbf{y}_{d}(t-w\delta t)\big),$ (5)
where the learned ROMs account for $w$ delayed latent points $\mathbf{y}_{d}(t-j\delta t)$ for $j=1,\ldots,w$. The latent points are obtained from previous predictions of $F_{\delta t}$ resulting in a closed-loop/recursive forecasting operator requiring only the initial $j=1,\ldots,w$ embeddings $\mathcal{R}(\mathbf{u}(j\delta t))$. This formulation enables efficient on-demand reconstruction of the high-dimensional space, as $F_{\delta t}$ evolves in latent space and lifting is performed only when a high-dimensional reconstruction is demanded.
@@ 2.2.1 From discrete pedestrian positions to continuous density fields
The initial step in our framework involves constructing a continuous density field from the discrete positions of pedestrians. Given microscopic observations, obtained by agent-based simulators in the form of Eq. (1, 2) or experimental data, macroscopic continuous quantities are typically derived via local averaging at each point of the spatial domain where the crowd moves. Here, we start from $N$ pairs of pedestrian positions $\{\mathbf{x}_{i}(t)\}_{i=1}^{N}\subset\Omega$ in a rectangular domain $\Omega=[a,b]\times[c,d]\subset\mathbb{R}^{2}$, as obtained by the SFM microscopic simulator, detailed in A. We next employ KDE [38] to estimate the continuous density fields along points on the spatial grid $\mathbf{x}\in\Omega$, as:
@@
where $K_{h}$ is a smooth non-negative kernel function satisfying $\int_{\Omega}K(\mathbf{x})\,d\mathbf{x}=1$. Using a standard uniform spatial discretization $\delta x=(b-a)/n_{x}$, $\delta y=(d-c)/n_{y}$ with $n_{x}$$n_{y}$ grid points along the xx/$y$-axis of the domain $\Omega$, the resulting from Eq. (6) density fields are computed as $\rho(\mathbf{x},t)=[\rho((x_{1},y_{1}),t),\ldots,\rho((x_{n_{x}},y_{n_{y}}),t)]^{\top}\in\mathbb{R}^{n_{x}\times n_{y}}$, where $\rho((x_{i},y_{j}),t)$ corresponds to the density computed at the grid point $(x_{i},y_{j})$, for $i=1,\dots,n_{x}$, $j=1,\dots,n_{y}$. Then, the total mass in the domain is given by $S(t)=\sum_{i=1}^{n_{x}}\sum_{j=1}^{n_{y}}\rho((x_{i},y_{j}),t)\,\delta x\,\delta y$. The above procedure can be applied in multiple groups of pedestrians to derive (normalized) group macroscopic density fields. While KDE can be used for extracting continuous macroscopic fields, such as momentum and energy, from any microscopic discrete quantities, here we only construct density fields. Additional details for the discretization of the computational domain $\Omega$, the choice of kernel $K_{h}(\cdot)$, the consideration of obstacles in $\Omega$ (where a binary mask function is employed to consider zero density at the grid points of the obstacle), and the treatment of periodic boundary conditions are provided in B.
@@ 2.2.2 The Restriction and Lifting Operators
where $n_{c}=n_{x}\times n_{y}$. To construct the restriction operator, we apply POD [104] to a dataset consisting of macroscopic density “solutions” in time, obtained by microscopic simulations via Eq. (6), for different initial conditions; see C for details. Denoting the densities by $\rho^{(c)}(\mathbf{x},t_{0}+k\delta t)\equiv\rho^{(c)}(\mathbf{x},t_{k})\in\mathbb{R}^{n_{c}}$, where $\delta t$ is a microscopic inner simulation step/observation sampling time of an experiment/real data, $t_{0}$ is the initial time, and the superscript $(c)$ denotes the case $c=1,\ldots,C$ of the different initial conditions considered, the training dataset generated by microscopic runs is represented by the matrix $X\in\mathbb{R}^{n_{c}\times(C\cdot K)}$:
$X=\Bigl\{\rho^{(c)}(\mathbf{x},t_{k})\,:\,k=1,\ldots,K,\,c=1,\ldots,C\Bigr\}.$ (8)
@@
where $U\in\mathbb{R}^{n_{c}\times n_{c}}$ is the matrix collecting the left-singular vectors, $\Sigma\in\mathbb{R}^{n_{c}\times n_{t}}$ is the matrix collecting the singular values, $V\in\mathbb{R}^{n_{t}\times n_{t}}$ is the matrix collecting the right-singular vectors, $I_{n_{t}}$ is the $n_{t}$-dim. unity matrix and $\mathbf{1}_{n_{t}}$ is the $n_{t}$-dim. unit column vector. The latent space is spanned by the first $d$ left-singular vectors forming $U_{d}$. Thus, the projection of $X$ on the latent space provides the low-dimensional embedding:
$Y_{d}=U_{d}^{\top}\bar{X},\,\,Y_{d}\in\mathbb{R}^{d\times n_{t}}.$ (11)
@@
We next consider a more challenging case of two interacting pedestrian groups in the domain $\Omega$, a typical scenario in dense counterflow settings. As in the one-group case, we collect the density profiles $\rho^{(c)}_{l}(\mathbf{x},t_{k})\in\mathbb{R}^{n_{c}}$ of each group $l=1,2$ via Eq. (6), and represent them in the snapshot matrices:
$X^{(1)}=\Big[x_{1}^{(1)},\ldots,x_{n_{t}}^{(1)}\Big]\in\mathbb{R}^{n_{c}\times n_{t}},\quad X^{(2)}=\Big[x_{1}^{(2)},\ldots,x_{n_{t}}^{(2)}\Big]\in\mathbb{R}^{n_{c}\times n_{t}},$ (19)
with columns $x^{(l)}_{k}=\rho^{(c)}_{l}(\mathbf{x},t_{k})/S_{l}(t_{k})\in\mathbb{R}^{n_{c}}$ for $k=1,\ldots,n_{t}=C\cdot K$. Applying the single‑group POD (Proposition 1) separately to $X^{(1)}$ and $X^{(2)}$ yields reconstructions that preserve each group’s density, but ignore any statistical coupling between the groups. To retain mass‑preservation while also incorporating dependencies among groups, we construct enriched projection bases that combine the individual POD modes with cross‑covariance modes, as shown in the following proposition.
Proposition 2.
Let us assume the snapshot matrices $X^{(1)},\,X^{(2)}\in\mathbb{R}^{n_{c}\times n_{t}}$ with columns being normalized density fields of two interacting population groups. Furthermore, let $W\in\mathbb{R}^{n_{c}\times m}$ and $T\in\mathbb{R}^{n_{c}\times m}$ be the left and right singular vectors of the cross-covariance matrix $C^{(12)}=\frac{1}{n_{t}}\bar{X}^{(1)}\left(\bar{X}^{(2)}\right)^{\top}\in\mathbb{R}^{n_{c}\times n_{c}}$, and consider the augmented orthonormal projection bases:
$\tilde{U}^{(1)}=\begin{bmatrix}\frac{1}{\sqrt{n_{c}}}\mathbf{1}_{n_{c}}&U_{d_{1}}^{(1)}&\hat{W}\end{bmatrix}\in\mathbb{R}^{n_{c}\times(d_{1}+m+1)},\quad\tilde{U}^{(2)}=\begin{bmatrix}\frac{1}{\sqrt{n_{c}}}\mathbf{1}_{n_{c}}&U_{d_{2}}^{(2)}&\hat{T}\end{bmatrix}\in\mathbb{R}^{n_{c}\times(d_{2}+m+1)},$ (20)
where $U_{d_{1}}^{(1)}\in\mathbb{R}^{n_{c}\times d_{1}}$ and $U_{d_{2}}^{(2)}\in\mathbb{R}^{n_{c}\times d_{2}}$ are the first $d_{1}$ and $d_{2}$ left singular vectors from the SVD of $\bar{X}^{(1)}$ and $\bar{X}^{(2)}$, respectively, and the terms $\hat{W}=W_{\perp}(W_{\perp}^{\top}W_{\perp})^{-1/2}$ and $\hat{T}=T_{\perp}(T_{\perp}^{\top}T_{\perp})^{-1/2}$ are computed by the orthogonal projections $W_{\perp}=\left(I_{n_{c}}-\mathbf{1}_{n_{c}}\mathbf{1}^{\top}_{n_{c}}/n_{c}-U_{d_{1}}^{(1)}(U_{d_{1}}^{(1)})^{\top}\right)W\in\mathbb{R}^{n_{c}\times m}$ and $T_{\perp}=\left(I_{n_{c}}-\mathbf{1}_{n_{c}}\mathbf{1}^{\top}_{n_{c}}/n_{c}-U_{d_{2}}^{(2)}(U_{d_{2}}^{(2)})^{\top}\right)T\in\mathbb{R}^{n_{c}\times m}$. If the original densities are mass-preserving along time steps, i.e., if $\mathbf{1}^{\top}_{n_{c}}X^{(1)}=\mathbf{1}^{\top}_{n_{t}}$ and $\mathbf{1}^{\top}_{n_{c}}X^{(2)}=\mathbf{1}^{\top}_{n_{t}}$, then the reconstructed densities computed by projecting $X^{(1)}$ and $X^{(2)}$ onto the augmented bases as:
$\tilde{X}^{(1)}=\tilde{U}^{(1)}\left(\tilde{U}^{(1)}\right)^{\top}\bar{X}^{(1)}+X^{(1)}(I_{n_{t}}-H),\quad\tilde{X}^{(2)}=\tilde{U}^{(2)}\left(\tilde{U}^{(2)}\right)^{\top}\bar{X}^{(2)}+X^{(2)}(I_{n_{t}}-H),$ (21)
are also mass-preserving, i.e., $\mathbf{1}^{\top}_{n_{c}}\tilde{X}^{(1)}=\mathbf{1}^{\top}_{n_{t}}$ and $\mathbf{1}^{\top}_{n_{c}}\tilde{X}^{(2)}=\mathbf{1}^{\top}_{n_{t}}$.
Proof.
Assume the original densities are mass-preserving for each group. From the properties of the single-population POD (see Eq. (15) and Eq. (16)), it follows that:
$\mathbf{1}_{n_{c}}^{\top}\bar{X}^{(l)}=\boldsymbol{0}^{\top}_{n_{t}},\quad\mathbf{1}_{n_{c}}^{\top}U^{(l)}_{d_{l}}=\mathbf{0}_{d_{l}}^{\top},$ (22)
where $\bar{X}^{(l)}=X^{(l)}H$ is the centered data matrix and $U_{d_{l}}^{(l)}\in\mathbb{R}^{n_{c}\times d_{l}}$ are the first $d_{l}$ left singular vectors from the SVD of $\bar{X}^{(l)}$ for the $l=1,2$ groups. Let’s consider now the SVD of the cross-covariance matrix:
$C^{(12)}=\frac{1}{n_{t}}\bar{X}^{(1)}\left(\bar{X}^{(2)}\right)^{\top}=W\Sigma_{C^{(12)}}T^{\top},$ (23)
with $W\in\mathbb{R}^{n_{c}\times m}$, $\Sigma_{C^{(12)}}\in\mathbb{R}^{m\times m}$ and $T\in\mathbb{R}^{n_{c}\times m}$, where $m$ denotes the number of retained cross-covariance modes. Each column of $C^{(12)}$ is a linear combination of the columns of $\bar{X}^{(1)}$, and each row a linear combination of the rows of $\bar{X}^{(2)}$. Consequently:
$\mathcal{R}(W)\equiv\mathcal{R}\big(C^{(12)}\big)\subseteq\mathcal{R}(\bar{X}^{(1)}),\quad\mathcal{R}(T)\equiv\mathcal{R}\big((C^{(12)})^{\top}\big)\subseteq\mathcal{R}\big(\bar{X}^{(2)}\big).$ (24)
To obtain interaction modes orthogonal to both the unity vectors $\mathbf{1}_{n_{c}}$ and the individual POD bases $U^{(l)}_{d_{l}}$ for $l=1,2$, we project the $W$ and $T$ onto the complementary subspace:
$W_{\perp}=\left(I_{n_{c}}-\frac{1}{n_{c}}\mathbf{1}_{n_{c}}\mathbf{1}^{\top}_{n_{c}}-U_{d_{1}}^{(1)}(U_{d_{1}}^{(1)})^{\top}\right)W\in\mathbb{R}^{n_{c}\times m},\quad T_{\perp}=\left(I_{n_{c}}-\frac{1}{n_{c}}\mathbf{1}_{n_{c}}\mathbf{1}^{\top}_{n_{c}}-U_{d_{2}}^{(2)}(U_{d_{2}}^{(2)})^{\top}\right)T\in\mathbb{R}^{n_{c}\times m},$ (25)
and define their orthonormalized versions as:
$\hat{W}=W_{\perp}(W_{\perp}^{\top}W_{\perp})^{-1/2},\quad\hat{T}=T_{\perp}(T_{\perp}^{\top}T_{\perp})^{-1/2}.$ (26)
By construction of $\hat{W}$ and $\hat{T}$, we obtain:
$\mathbf{1}_{n_{c}}^{\top}\hat{W}=\mathbf{0}^{\top}_{m},\quad(U_{d_{1}}^{(1)})^{\top}\hat{W}=0\in\mathbb{R}^{d_{1}\times n_{c}},\quad\mathbf{1}_{n_{c}}^{\top}\hat{T}=\mathbf{0}^{\top}_{k},\quad(U_{d_{2}}^{(2)})^{\top}\hat{T}=0\in\mathbb{R}^{d_{2}\times n_{c}}.$ (27)
The augmented orthonormal bases (Eq. (20)) are therefore:
$\tilde{U}^{(1)}=\begin{bmatrix}\frac{1}{\sqrt{n_{c}}}\mathbf{1}_{n_{c}}&U_{d_{1}}^{(1)}&\hat{W}\end{bmatrix}\in\mathbb{R}^{n_{c}\times(d_{1}+m+1)},\quad\tilde{U}^{(2)}=\begin{bmatrix}\frac{1}{\sqrt{n_{c}}}\mathbf{1}_{n_{c}}&U_{d_{2}}^{(2)}&\hat{T}\end{bmatrix}\in\mathbb{R}^{n_{c}\times(d_{2}+m+1)},$ (28)
according to which, the reduced coordinates of the snapshots can be obtained as:
$Y^{(1)}=(\tilde{U}^{(1)})^{\top}\bar{X}^{(1)}\in\mathbb{R}^{(d_{1}+m+1)\times n_{t}},\qquad Y^{(2)}=(\tilde{U}^{(2)})^{\top}\bar{X}^{(2)}\in\mathbb{R}^{(d_{2}+m+1)\times n_{t}}.$ (29)
Multiplying the augmented basis in Eq. (28) by $\mathbf{1}_{n_{c}}^{\top}$ from the left, and using Eq. (27) together with the orthogonality $\mathbf{1}_{n_{c}}^{\top}U^{(l)}_{d_{l}}=\mathbf{0}_{d_{l}}^{\top}$ in Eq. (22) gives:
$\mathbf{1}_{n_{c}}^{\top}\tilde{U}^{(1)}=\sqrt{n_{c}}\mathbf{e}_{1}^{\top},\quad\mathbf{1}_{n_{c}}^{\top}\tilde{U}^{(2)}=\sqrt{n_{c}}\mathbf{e}_{1}^{\top},$ (30)
where $\mathbf{e}_{1}=[1,0,\ldots,0]^{\top}$. Hence, for the projected centered data, we get:
$\mathbf{1}_{n_{c}}^{\top}\tilde{U}^{(l)}(\tilde{U}^{(l)})^{\top}\bar{X}^{(l)}=\sqrt{n_{c}}\mathbf{e}_{1}^{\top}(\tilde{U}^{(l)})^{\top}\bar{X}^{(l)}=\mathbf{1}_{n_{c}}^{\top}\bar{X}^{(l)}=\mathbf{0}_{n_{t}}^{\top},\quad l=1,2.$ (31)
The last equality uses the mass-preservation property for the original densities $\mathbf{1}_{n_{c}}^{\top}\bar{X}^{(l)}=\boldsymbol{0}^{\top}_{n_{t}}$ in Eq. (22).
Finally, using the reconstruction formula in Eq. (21), and the identities derived in Eqs. (22) and (31), we obtain:
$\mathbf{1}^{\top}_{n_{c}}\tilde{X}^{(l)}=\mathbf{1}_{n_{c}}^{\top}\tilde{U}^{(l)}(\tilde{U}^{(l)})^{\top}\bar{X}^{(l)}+\mathbf{1}^{\top}_{n_{c}}X^{(l)}(I_{n_{t}}-H)=\mathbf{1}^{\top}_{n_{c}}X^{(l)}(I_{n_{t}}-H)=\mathbf{1}_{n_{t}},\quad l=1,2.$ (32)
Thus, the reconstructions $\tilde{X}^{(1)}$ and $\tilde{X}^{(2)}$, obtained via the projection of $X^{(1)}$ and $X^{(2)}$ onto the augmented basis in Eq. (28), retain the mass-preservation property of the original data. ∎
@@
In summary, given a density distribution $\rho(\mathbf{x},t)$ obtained via Eq. (6) for a single population, we obtain the normalized density along the $n_{c}$ grid points of $\Omega$ as:
$x(t)=\dfrac{1}{S(t)}\big[\rho((x_{1},y_{1}),t),\ldots,\rho((x_{n_{x}},y_{n_{y}}),t)\big]^{\top}\in\mathbb{R}^{n_{c}},$ (33)